### 14 Feb 2024 v1

# Structured Language Generation Model for Robust Structure Prediction

Abstract
We propose Structured Language Generation Model (SLGM), a mixture of new loss function and inference method for better generalization of structured outputs. Previous studies on structure prediction (e.g. NER, RE) make use of explicit dataset information, which would boost performance, yet it might pose challenges to robust generalization in real-world situations. Instead, our model gives generalized format information about data indirectly. With format information, we could reduce sequence-to-sequence problem into classification problem via loss calibration and formatted decoding. Our experimental results showed SLGM successfully maintain performance without dataset information, and showed much less format errors. We also showed our model can work like adapters on individual dataset, with no additional training.
1 Introduction
Recent advances in pre-trained language models (PLMs) and large language models (LLMs) allow computational approaches to more tasks and problems than ever. Although they are not yet capable of human-like language learning and understanding, their general purpose and end-to-end capabilities remain an attractive option in solving natural language understanding (NLU) tasks.
Structure prediction is among many attempts to solve multiple NLU tasks with a single model. Structure prediction is set of natural language understanding tasks, including named entity recognition, relation extraction, and other tasks that requires structure understanding. To perform structure prediction, it is vital that a language model to have a thorough understanding of language and its structure. However, some studies have demonstrated that LLM themselves lack the capability to understand language structure effectively Zhong et al. (2023); Berglund et al. (2023). We observed that ChatGPT-4111https://chat.openai.com  (OpenAI, 2024), even when equipped with detailed instructions and examples, fall short comparing with relatively smaller models, as seen in Table 1. To overcome such shortcomings, some studies like DeepStruct (Wang et al., 2023) and UIE (Lu et al., 2022) proposed pre-training on structured dataset, which contains large number of entities and relations, to improve generative models’ understanding ability. DeepStruct used large corpuses like TEKGEN and KELM  (Agarwal et al., 2021), which is based on WikiData (Vrandečić and Krötzsch, 2014). UIE used Wikipedia and ConceptNet  (Speer et al., 2018) as their pre-training corpus.
Previous works on structure prediction like TANL (Paolini et al., 2021) and Deepstruct reported performance improvements in various structure-related tasks with single model. We believe the improvements represent dataset-specific engineering, however, rather than improvement in structure understanding, as the improved models are fine-tuned or given explicit dataset information. Performing inference without explicit dataset information supports this, as we observe a significant drop in benchmark scores and witness a diverse array of formatting errors, as shown in Table  1 and Section  4. We aim to measure and optimize on general structure understanding abilities, rather than task-specific engineering abilities. We also focus on robustness, where the model performs consistently well in a diverse set of situations, including low-resource settings.
Model Ent. F1 Rel. F1
DeepStruct Original 88.4 72.8
- Dataset 81.7 40.7
+ Tagset
- Dataset 79.6 48.2
TANL Original 90.3 70.0
- Dataset 24.0 2.8
ChatGPT4* 57.7 34.1
Table 1: F1 Score on CoNLL-04 joint entity and relation dataset of previous works, regarding to the inclusion of dataset or tagset information, and ChatGPT4. We used DeepStruct 10B multi-task model and TANL multi-task model for inference. Score of ChatGPT is measured on first 40 sentence of test dataset. Full prompt we used to inference on ChatGPT4 is shown in Appendix  A
Figure 1: Difference between input structure of previous works and SLGM on NER task. Blue area is task specific information. Green area is dataset specific information. Instead of using dataset name as prompt, we augmented additional format information into model.
In this paper, we propose a new framework named Structured Language Generation Model (SLGM) which can maintain robust structural understanding ability. Our framework comprise of task-specific format information, format loss, and formatted inference. The SLGM framework is illustrated in Figure 1. Task-specific information contains structure of desired outputs with tagset. Format loss works to learn the intended format, and formatted decoding encourages the model to decode with format guidelines.
We trained and evaluated our model on structure prediction tasks, demonstrating significant performance improvements over models of the same size that were trained on base sequence-to-sequence task which utilizes only cross entropy. Also, our model successfully replaced dataset information. Despite our framework not undergoing actual training, it follows performance of dataset-specific engineered models.
The contributions of this work are as follows.
  • We introduce SLGM, which maintain structural understanding ability well without dataset information. (Section  3)
  • We trained and evaluated SLGM on structured prediction task, and observed improvements on almost every task compared to plain cross entropy loss. (Section  4,  5)
  • With additional studies, we showed our method can work as zero-weight adapter that simulates dataset specific fine-tuning, and works better on low-resource environments. (Section  6)
2 Related Works
Since emergence of Attention mechanism  (Vaswani et al., 2017), many pre-trained language models were introduced to solve NLP tasks. Models like T5 (Raffel et al., 2023), Bart (Lewis et al., 2020), Flan-T5 (Chung et al., 2022), GLM (Du et al., 2022) tried to transfer NLP tasks into sequential generation problem. These models introduced new pre-training methods to solve general tasks, and they showed improvement on zero-shot task scores.
Structure prediction task contains many NLU tasks like NER, RE, SRL. It is part of multi-task problem, and requires model to understand structure of languages to achieve high score. Traditional approach, however, created task-specific models for each task. Based on sequence to sequence pretrained LMs, studies like TANL (Paolini et al., 2021), DeepEx (Wang et al., 2021), DeepStruct (Wang et al., 2023), UIE (Lu et al., 2022), USL (Min et al., 2024) created unified model to solve structure understanding tasks in single model. TANL and UIE created specific output formats for each task. DeepEx and DeepStruct translated every output as triples to generate more unified output across tasks. Additionally, DeepStruct, UIE and USL used structural pre-training method with additional data to improve model’s structural understanding ability. Our model adopted some idea from generation tasks to replace dataset specific information or additional training.
In sequence generation tasks, there has been many research on methodologies for integrating words from the original text into generation result. Copying mechanism  (Gu et al., 2016) introduced a new method computing the generation score of a token alongside the score of an extracted token from the original text, which enables the direct inclusion of tokens from the original source. This approach was broadly adopted, from summarization  (Zeng et al., 2016) to dialogue systems  (Park et al., 2023).  (Choi et al., 2021) extended copy mechanism, by explicitly decide when to copy and when to generate. By adjusting the weights, it can also be utilized for classification of tokens within a sentence.
Not only using source on generation, there were many attempts using formats or masks when generating sequence. Before machine learning approach were general in NLP area, there were many rule-based approaches (Hovy, 1993). Since the widespread adoption of Pretrained Language Models (PLMs), however, the use of formats or rules for generation has not been extensively researched. On abstract summarization,  (Gehrmann et al., 2018) masked some part of source document, not to copy from useless parts. On machine translation (Ghazvininejad et al., 2019) to speed up inference, or sentiment analysis (Li et al., 2020) to create more precise augmentation data. Among generation tasks, code generation (Ling et al., 2016) is one of the most sensitive study related to format. Study like CodePAD (Dong et al., 2023) utilized pushdown automata to generate tokens according to current state. Comparing with code generation, structure prediction has less states and simple, concrete state transition rules, which can be implemented easily.
We thought structured prediction task can be reduced into some point between extractive summarization and classification, with much tighter requirements. We adopted concept from formatted decoding of code generation and copy mechanism of summarization to get higher accuracy and less format misses on structure prediction.
3 Method
Our model operates around a task and data-specific format. We have established a tagset and a specific set of words related to the task and dataset, in order to guide the model about what it should generate. SLGM uses this format to calculate format loss, and formatted decoding.
3.1 Task-specific Format
Tasks like named entity recognition and relation extraction need to extract its element from source sentence. Moreover, within the same task, the tagset can vary depending on the dataset. Previous works inferred valid tagset from dataset names. We introduce additional task-specific format information, which also contains valid tagset, and convey this to model distinct from the input prompt. Task-specific format is used for calculating additional loss and masked inference.
Task-specific format is defined with separator tokens and area between separators, which is called Slots. A single format string contains multiple separator tokens (<;>) to distinguish slots and a triple separator token (</>) to indicate the end of the format. Each slot can contain a specific string or either two special slot tokens - <ANY> and <SOURCE>. When <ANY> token is given at the slot, it means the model can generate any token. <SOURCE> token means the model should generate tokens only appeared in source sentence. If specific string is given in slot, it means the model should generate one of those tokens given in current slot, which can be used as tagset information. During input generation process, we parse and tokenize the given format string into a format mask tensor. Format mask tensor is a boolean tensor that has true value on legal token ids per each slot. This tensor is used on calculating format losses or masking illegal tokens on decoding process.
3.2 Format Loss
Traditional cross entropy loss used in previous works is calculated over every possible vocabulary defined in model. We thought this is way too broad, so in addition to cross entropy, we introduce structure loss for better format prediction, and slot loss for better tagset prediction.
3.2.1 Structure Loss
Structure loss aims to produce outputs in correct format. We fixed every output into triple format, so it is important to correctly produce separator tokens.
$L_{st}=\sum_{t\in S}l_{t}\cdot missed\cdot w_{miss}$
$l_{t}=-log\frac{e^{x_{y_{t}}}}{\sum_{c\in V}e^{x_{t}}}$
where $S$ is location of separator tokens, $w_{miss}$ is miss weight, which is hyperparameter. $missed$ refers to the number of instances where a separator did not have the highest generation probability at its designated position within a sentence. Since wrong generation of separator gives critical parsing issue, structure loss gives bigger penalty of wrong separator prediction.
3.2.2 Slot Loss
Slot loss makes model produce correct tokens among filtered tokens. For example, NER tasks should select tokens from original sentence for head, and specific tags defined in dataset for tail. The formula for calculating slot loss resembles that of cross entropy; however, the key distinction is in the denominator’s token size. Slot loss computation only involves a pre-selected set of legal tokens pertinent to the slot, rather than encompassing all tokens. With slot loss, we can translate sequence generation task into classification task, which has far smaller classes than whole vocabulary size. Slot loss is defined as:
$L_{sl}=\sum_{k=0}^{N}l_{t,k}(t\in S_{k})$
$l_{t,k}=-log\frac{e^{x_{y_{t}}}}{\sum_{c\in V_{k}}e^{x_{t,c}}}$
where $N$ is number of slots, $S_{k}$ is slot $k$, $V_{k}$ is legal token set defined at slot $k$.
Final training loss is calculated as weighted sum of three losses:
$L=w_{ce}L_{ce}+w_{st}L_{st}+w_{sl}L_{sl}$
where $ce$ stands for cross entropy, $st$ for structure, and $sl$ for slot. $w$s are loss weight, which are hyperparameters.
3.3 Formatted Decoding
During inference time, we hold state information of current generation per each sentence. This state information works as simple finite state machine. Each time when the model generates slot separator, state counter increases. When model generates triple separator, state counter sets to 0. Before generating the next token, we retrieve the format mask corresponding to the current generation state of the slot, enforcing the generation of legal tokens by adding a large negative value to the scores of illegal tokens.
Task Dataset Format
NER CoNLL-03  (Sang and De Meulder, 2003) <SOURCE> <;> instance of <;> tagset </>
Ontonotes-v5  (Weischedel et al., 2013)
GENIA  (Kim et al., 2003)
CoNLL-04  (Roth and Yih, 2004)
NYT  (Riedel et al., 2010)
RE TACRED  (Zhang et al., 2017) <SOURCE> <;> tagset <;> <SOURCE> </>
TACREV  (Alt et al., 2020)
CoNLL-04  (Roth and Yih, 2004)
NYT  (Riedel et al., 2010)
SRL CoNLL-12  (Pradhan et al., 2012) <SOURCE> <;> instance of <;> tagset </>
ID ATIS  (Hemphill et al., 1990) intent <;> is <;> tagset </>
SNIPS  (Coucke et al., 2018)
DST MultiWOZ  (Budzianowski et al., 2018) [User] <;> <SOURCE> <;> <ANY> </>
Table 2: Task, dataset and task-specific formats. Joint entity and relation extraction corpus like CoNLL 2004 and NYT dataset split one sentence into two sentences, with different prompt and format.
4 Experiments
Our research followed the DeepStruct (Wang et al., 2023) experiment protocol, involving two stages: Structural pre-training and multi-task training. DeepStruct also fine-tuned multi-task model with each dataset to get higher result. However, since our main goal is to enhance model performance in multi-task scenarios, we report our main score on multi-task training.
4.1 Tasks and Datasets
4.1.1 Structural Pre-training
We used TEKGEN and KELM  (Agarwal et al., 2021) as our pre-training corpus. We extracted entities and relations from given triples, and generated named entity recognition and relation extraction corpus. Since TEKGEN and KELM only contains surface form of entity but nothing about entity type, we retrieved entity type using WikiData (Vrandečić and Krötzsch, 2014). There are about 50k wikidata entity types and 40k relation types, so instead of using them directly, we mapped frequent entity and relation types into common types. Detailed statistics of mapping result is shown in Appendix  B. We used 100k sentences from each corpus, NER and RE respectively (i.e. total 400k sentences).
4.1.2 Multi-task training
For multi-task training, we used dataset shown on Table  2 with given format for multi-task training. Following TANL and DeepStruct, we marked brackets around targets when needed, like head and tail of TACRED relation extraction dataset, or predicates of CoNLL-12 semantic role labeling dataset. We trained 2 epochs using every dataset on structural pretrained model.
4.2 Baseline
We compared our model with various training settings that have same structure, same size, and same data. First, we trained model using cross entropy loss, using same input prompt and answer with SLGM(CE). Also, we trained model using DeepStruct’s prompt (shown in Figure  1), which includes task and dataset information, with cross entropy loss as our oracle model (CE+Data). To see how dataset information affects to multi-task setting, we additionally evaluated DeepStruct prompt model without dataset information(CE+Data*).
4.3 Evaluation Method
We report score and format errors on each dataset. On every task except dialogue state tracking, we evaluated with micro F1 score. For joint entity and relation extraction, we recorded F1 score of entity and relation respectively. For dialogue state tracking task, we used joint accuracy score. We categorized format errors into three types: length mismatch, source mismatch, and tagset mismatch. Length mismatch means length of generated tuple does not match with given format. Source mismatch means the element of output violated <SOURCE> slot token restriction. When the model generated tokens that don’t exist in source sentence, source mismatch count increases. Finally, a tagset mismatch occurs when the model generates an output that is not defined within the tags specified by the dataset.
Task Dataset CE CE+Data* SLGM CE+Data
Score FE Score FE Score FE Score FE
NER CoNLL-03 64.03 5843 10.78 9832 84.18 25 89.79 15
OntoNotes 71.89 717 24.88 4525 81.64 99 83.84 83
GENIA 0.00 1137 66.42 26 72.24 7 67.37 35
RE TACREV 54.37 5 67.99 10 72.93 2 66.72 7
TACRED 49.18 5 60.32 10 64.32 2 57.70 7
JER CoNLL-04 0.00 794 0.00 768 71.17 0 74.43 11
4.87 530 34.24 442 33.90 2 46.43 99
NYT 85.35 475 78.60 214 89.80 95 88.83 37
54.79 82 40.85 2346 64.75 11 66.01 38
SRL CoNLL-12 80.22 20 82.31 18 84.72 0 82.15 16
ID ATIS 92.64 31 93.62 13 96.52 0 94.36 10
SNIPS 95.22 1 96.01 0 96.72 0 96.86 0
DST MultiWOZ 37.75 10 35.03 42 40.96 0 39.33 4
Average 53.10 742.31 53.16 1403.54 73.37 18.69 73.37 27.85
Table 3: Main result of our experiment on multi-task setting. FE stands for "Format Error". Scores are higher the better, and format errors are lower the better. Average scores and format errors are dataset-wise macro average.
4.4 Implementation Detail
We used Flan-T5 (Chung et al., 2022), an instruction-tuned version of T5 from HuggingFace hub  (Wolf et al., 2020) as our base model. As shown in Figure  1, we used "Extract target from: sentence" as our input prompt, where target is determined by task. We trained our model using base sized checkpoint(220M parameters), with batch size 16 per GPU on 4 NVIDIA A100 GPUs with 40G memory. For SLGM hyperparameters, we used $w_{ce}=$ 0.5, $w_{st}=$ 0.2, $w_{miss}=$ 0.33, and $w_{sl}=$ 0.3. We used greedy decoding when generating answers.
5 Main Result
5.1 Comparing with plain extraction
Table  3 shows our full result. Our model showed better result than models without dataset information, and showed comparative result with dataset-aware model. Also, our model generated less format errors than dataset-aware model, which mean our model is much robust on real world dataset. When training without dataset information, we observe an average performance drop of 10 points compared to when dataset information is included. Furthermore, the frequency of format errors increases almost tenfold. Training with dataset information and then removing it yields worse results compared to training without the dataset information from the start. This means model mostly depends on dataset information.
However, it’s important to note that CE and CE+Data* operate without any tagset information, whereas SLGM can indirectly acquire tagset information. Since there are multiple datasets for NER and JER, which have their unique tagset, the occurrence of format errors is inevitable. Looking for RE and ID, which share identical or similar tagsets, and for SRL and DST, where only one dataset exists, it is observed that the baseline model also exhibits a low number of format errors. In the following section, we further analyze format errors and experiment to determine whether the use of formatted decoding as a means of providing indirect tagset information is effective for the baseline model as well.
5.2 Comparing with models using formatted decoding
To investigate the impact of formatted decoding on generating the correct format, we applied formatted decoding stage to baseline models and, conversely, removed from our SLGM model before proceeding with inference. Table  4 shows the average score and average format errors across all datasets. Formatted decoding itself can work greatly on extracting correct tagsets. On CE model trained without dataset information, model using formatted decoding achieved 16 point higher F1 score and 5% format errors compared with original model. However, when trained with dataset information, the model cannot recover from absence of dataset information even with format information. F1 score and format errors are improved, but still showed large number of format errors comparing with other models using formatted decoding. When designing models with real-world applications in mind, it becomes evident that including dataset names during training can lead to undesirable effects.
When using format loss, on the other hand, showed best score and had least format errors among models which did not use formatted decoding. We think this means format loss itself, works as dataset indicator. Slot losses produces tagset-specific loss, and this applies stronger back-propagation than cross entropy, which applies on every tokens. This makes model know more about what to extract on given sentence, which indicates the model understands structure of sentence better.
Full result and detailed analysis on each dataset is shown on Appendix  C.1
Loss FD Avg. F1 Avg. FE
CE
53.10 742.31
O 69.87 37.46
CE+Data*
53.15 1403.53
O 59.03 548.23
CE+FL
66.74 503.76
O 73.37 18.69
Table 4: Average score regarding to loss and inference method. FD stands for Formatted Decoding. FL on last two rows stands for Format Loss. Note that model trained with cross entropy and format loss, using formatted decoding(CE+FL+FD) is equal to SLGM.
5.3 Format Error Analysis
We analyzed ratio of format errors for each model, with and without formatted decoding. Figure  2 shows format error count and error type ratio across tasks. As illustrated in the figure, the absence of dataset information most prominently results in tagset mismatches. Formatted decoding can reduce tagset mismatch dramatically. It could not completely remove them, since formatted decoding only masks on tokens, not sequences. In cases where tagset mismatches were not resolved, the majority involved generating shorter tags with the same meaning, such as producing only "LOC" when "LOCATION" was required. For reference model trained with dataset information, source mismatch were more frequent than other models. Since it is tuned with dataset information, the model can clearly distinguish legal tagset. However, it cannot correctly infer correct source tokens, especially in named entity recognition.
Figure 2: Total format error counts and ratio per tasks. Empty area means there are no format errors in such setting. Number in circle means sum of format errors across every dataset inside given task.
6 Additional Studies
We conducted additional studies to see in what situation our model works well. For each subsection, we report average result across all datasets. Full result is shown on Appendix  C.
6.1 Fine-tuning
Loss FT Avg. F1 Avg. FE
CE
53.10 742.31
O 77.70 56.45
CE+Data*
53.15 1403.53
O 63.53 606.00
CE+FL
66.74 503.76
O 78.93 16.09
SLGM
73.37 18.69
O 78.85 0.64
Table 5: Average score and format error regarding to fine-tuning. FT stands for fine-tuning. SLGM equals to CE + Format loss + Formatted decoding.
We sought if our model is competitive enough with fine-tuning. Since fine-tuning involves training on a single dataset with clean tagset, it can be expected to improve performance even for models without dataset information. Based on models trained in multi-task environment, we conducted an additional 5 epochs of training for each dataset and then evaluated their performance. Table  5 shows the result. Comparing with result of Table  4, fine-tuning can achieve relatively higher F1 score than formatted decoding. Furthermore, when looking at fine-tuning result of CE+FL and SLGM, we can observe that formatted decoding has minimal impact on performance. These results lead us to believe that formatted decoding can partially replicate the effects provided by fine-tuning. This can be seen as similar to the principle of operation of adapters  (Houlsby et al., 2019), in that the original model remains unchanged, and additional lightweight features are incorporated.
6.2 Model size
Figure 3: Average score and format error according to model size(small-base-large). Solid line indicates Average F1 score(left side). Dotted line indicates format errors(right side).
We additionally evaluated our loss function and formatted inference on Flan-T5-small(77M) model and Flan-T5-large(783M) model. Train and inference setting is same as written in Section  4. Summary of results is shown in Figure  3.
Our model also showed better result than CE with dataset information on small sized model. Since smaller models that doesn’t have enough language capacity are sensitive on additional information, our format loss and formatted inference provides stronger hint on correct labels than plain cross entropy loss. On large sized model, plain cross entropy model got much better result than smaller models. With more model power, it can recognize characteristics of each dataset without any hints. Giving indirect hint with format loss, however, showed negative effect on large size. Still, with formatted decoding, model with format loss performed better than cross entropy model. We can infer format loss and formatted decoding should mixed together to get better results.
When using our best methods, we achieved state of the art performance in some datasets (Table  C.3). We achieved 94.8 micro F1 score on CoNLL-03 NER dataset, and 98.3 micro F1 score on ATIS intent detection dataset, which are both state of the art. This indicates that our method can further improve through fine-tuning, and it can achieve better performance without the need for engineering specific to dataset features.
6.3 Low-resource setting
Figure 4: Average score and format errors according to dataset fraction. Solid line represents model with structural pre-training, and dotted line represents model without structural pre-training.
We hypothesized that SLGM effectively retains the characteristics of data, prompting us to investigate its performance in low-resource environments through experimentation. For each multi-task dataset, we extracted data with proportional tag distributions, with at least one example, to use as training data. We didn’t used MultiWOZ dataset since it is hard to distinguish unique tags. To compare true power of our model, we trained in two different environment, with and without structural pre-training.
Figure  4 shows our experiment result. Regardless of structural pre-training, SLGM worked better than other models. SLGM even showed higher score and lower format errors than dataset aware model on low-resource environment. We also observed that formatted decoding is better to use than format loss on low-resource setting. Additionally, it is notable that models using format loss showed worse score when trained without structural pre-training and have more data.(CE+FL, SLGM) In our observation, the model showed stable scores across most dataset, but showed steep increase of hallucinations in NYT relation extraction task. We speculate that the use of format loss in low-resource environments makes it challenging to accurately learn the proper termination point, resulting in repetitive outputs.
Comparing dotted line with solid lines, structural pre-training has great improvements on all training settings. With structural pre-training, models did not exhibit a decline in performance regardless of dataset fraction. This shows structural pre-training enhances the stability of the model and can be particularly helpful in guiding how the model interprets sentences, especially when data is scarce.
7 Conclusion and Discussion
We introduced a new framework named SLGM, which utilizes task-specific formats. By using formats with format loss and formatted decoding, we reduced structure prediction task into classification tasks. We showed our models were more robust on multi-task environments without dataset information, which is more powerful on real-world scenario. Our method worked similarly as using fine-tuning on each dataset. With using bigger model and fine-tuning, our model still works better and achieved state of the art on some datasets.
We successfully replaced dataset information into formats, which can be expanded in various ways. Our research has only applied the fundamental concepts, and we think formats and formatted decoding can be more generalized. For example, we can remove unnecessary fixed slots, or create more complex formats with expanding decoding states using finite state machine. Our model has possibility on those situations too. We anticipate that format loss and formatted decoding will be effectively utilized in fields beyond structured prediction.
Limitations
Format loss and formatted decoding represent basic strategies designed to compel the model to produce outputs in accordance with a predetermined format. There may be many improvements on both scoring and real-world usage. There may be many problems in specific situations. Suppose we extract named entity with location but there is no location tag in format. The model expected location and extracted some words, but the location tag is illegal. In plain decoding strategy, the model would generate any token similar to location. However, when using formatted decoding, it would generate unexpected string, which could be worse than extracting wrong tag.
This scenario could potentially be addressed by employing a beam decoding strategy, which operates on triple units. This issue is related to the fact that the model does not know about format during actual attention calculation. Despite explicit format is given, model cannot utilize format information at attention layers. We think additional cross attention layer over formats may show better results. We leave these possible improvements as future works.
Ethics Statement
In our research, we utilized a pretrained model and made use of publicly available datasets published on the web. We acknowledge that the data obtained from the web may contain potential biases. We must clarify that our research did not specifically address or consider these biases. We ensured that all datasets employed in our study were accessed and used in a manner that respects their intended use and complies with any associated licenses or terms of service. We are also mindful of the potential biases present in these datasets and the pretrained model, and have endeavored to address these concerns in our research methodology.
We used ChatGPT as a tool for generating and refining text content. We acknowledge and address the ethical considerations associated with the use of such AI technology, and have thoroughly reviewed the content to ensure that it does not include any unethical material.
Appendix A Instruction of ChatGPT4
The result of Table  1 is measured with custom ChatGPT 4. We used custom prompt which defines task and tagset, with two examples to show exact format. Full prompt is shown below:
You are given multiple sentences. You have to extract relations between entities in given sentences. For each sentence, extract every named entity that have relations first. Types of named entity must be one of "location", "organization", "other", "human". Then, extract relations between extracted named entities. Types of relation must be one of "kills", "lives in", "works for", "located in", and "organization based in".
You must FOLLOW given structure. When extracting named entities, you must extract in this form: [ENTITY] <;> instance of <;> [TYPE] . When there are multiple named entitis. separate with </> . When extracting relations, you must extract in this form: [HEAD] <;> [TYPE] <;> [TAIL] . Same with named entity extraction, if there are multiple relations, separate with </> .
For each sentence, first line should be result of named entity extraction, and second line should be result of relation extraction. When given multiple sentence(separated by empty line), pad empty line between extractions.
When given a file, each line contains single sentence.
Example
Input
John Wilkes Booth , who assassinated President Lincoln , was an actor .
The opera company performed at the Palace of Fine Arts , in San Francisco , on June 30 and July 1-2 , said Kevin O ’Brien , a spokesman for the theater.
<input end>
Output
John Wilkes Booth <;> instance of <;> human </> President Lincoln <;> instance of <;> human
John Wilkes Booth <;> kills <;> President Lincoln
Palace of Fine Arts <;> instance of <;> location </> San Francisco <;> instance of <;> location </> June 30 <;> instance of <;> other </> July 1-2 , <;> instance of <;> other </> Kevin O ’Brien <;> instance of <;> human
Palace of Fine Arts <;> located in <;> San Francisco
<output end>
With given prompt, we uploaded test file and got result of first 40 sentences.
Appendix B Type Mapping of Pre-training Data
We collected entity types appeared more than 20k times, and manually labeled them into 5 different entity types: Person, Location, Organization, Product, and Event. Surface forms that we could not find from WikiData are not used. Table  6 shows example and part of statistics of entity type mapping. Table  7 contains full statistics of mapped entity types.
Original type Count Mapping Example
Human
9.7M Person
Umberto II of Italy
Country
4.2M Location
England
Taxon
1.2M -
Natuna Island surili
association football club
954k Organizaiton
FC Baden
Film
515k Product
Avengers
Summer Olympic Games
185k Event
2024 Summer Olympics
Table 6: Entity type mapping examples
Similarly, we collected relations containing entities that have valid types mapped at Table  6. Among these relations, we filtered out relations appeared less than 10k times. There were 168 relation phrases meeting these conditions. We additionally clustered 82 of 168 relation phrases into 35 similar clusters. Table  8 shows statistics of mapped relations.
Type Count
Location 10,594,011
Person 10,239,553
Organization 3,182,685
Product 2,233,553
Term 1,186,434
Event 514,083
Table 7: Entity type mapping statistics
Type Count
country 3,233,250
member of 1,768,459
located in 1,657,574
product of 1,034,479
date of birth 1,009,859
occupation 937,797
instance of 879,955
place of birth 737,735
educated at 496,122
cast member 446,835
works at 353,896
date of death 300,792
awarded 299,818
contains 253,338
distributor 162,945
part of 146,352
gender 142,878
place of death 122,463
start date 110,483
sibling 92,805
child 84,473
parent 81,600
owner 77,977
genre 76,344
spouse 76,093
nominated 69,833
winner 61,688
position 60,142
created 49,232
background 42,186
capital 39,123
twin towns 38,671
capital of 31,669
president 20,402
family 15,840
Table 8: Relation type mapping
Appendix C Detailed experiment result
This section shows full result of Section  5 and  6. We report scores and format errors on every dataset for each settings.
C.1 Formatted decoding
Task Dataset CE CE+FD CE+FL SLGM CE+Data*+FD CE+Data+FD
Score FE Score FE Score FE Score FE Score FE Score FE
NER CoNLL-03 64.03 5843 80.36 43 72.76 4437 84.18 25 26.80 2509 89.74 0
OntoNotes 71.89 717 78.20 191 81.30 411 81.64 99 25.36 4566 83.74 4
GENIA 0.00 1137 66.19 15 72.37 31 72.24 7 66.29 0 67.18 0
RE TACREV 54.37 5 65.21 2 72.97 7 72.93 2 67.98 3 66.74 2
TACRED 49.18 5 56.92 2 64.35 7 64.32 2 60.32 3 57.74 3
JER CoNLL-04 0.00 794 67.57 0 6.88 874 71.17 0 66.67 0 74.43 0
4.87 530 27.28 0 23.32 425 33.90 2 27.71 2 43.86 1
NYT 85.35 475 88.26 232 89.79 215 89.80 95 78.19 3 88.73 1
54.79 82 66.74 1 65.01 120 64.75 11 41.26 40 66.07 5
SRL CoNLL-12 80.22 20 82.51 1 84.72 14 84.72 0 82.30 1 82.15 1
ID ATIS 92.64 31 93.72 0 96.51 8 96.52 0 93.39 0 94.62 0
SNIPS 95.22 1 96.43 0 96.72 0 96.72 0 96.01 0 96.86 0
DST MultiWOZ 37.75 10 38.89 0 40.96 0 40.96 0 35.16 0 37.86 0
Average 53.10 742.31 69.87 37.46 66.74 503.77 73.37 18.69 59.03 548.23 73.05 1.31
Table 9: Full result of F1 scores and format errors regarding to formatted decoding.
Table  9 shows full result using formatted decoding (Section  5.2). Comparing with Table  3, formatted decoding greatly reduces format errors on NER and JER tasks. It also shows great score improvements on low resource data like CoNLL-04.
C.2 Fine-tuning
Task Dataset CE+FT CE+FD+FT CE+FL+FT SLGM+FT CE+Data*+FT CE+Data+FT
Score FE Score FE Score FE Score FE Score FE Score FE
NER CoNLL-03 91.63 10 92.29 0 93.65 4 93.65 0 77.99 2733 91.65 7
OntoNotes 82.26 376 82.23 4 84.22 11 84.21 0 41.93 2860 82.29 365
GENIA 75.78 5 68.09 0 75.44 6 75.42 0 66.72 18 75.70 4
RE TACREV 75.18 8 74.57 3 76.95 6 76.90 2 73.12 7 74.58 10
TACRED 65.78 8 64.94 3 66.87 6 66.82 2 64.96 7 65.51 10
JER CoNLL-04 83.50 2 82.68 0 85.60 3 85.48 0 0.00 877 84.86 1
52.43 6 50.85 0 55.71 14 55.24 0 36.35 126 50.36 4
NYT 90.11 11 90.08 0 90.66 13 90.68 0 85.28 15 90.36 20
74.70 35 74.47 0 76.78 102 76.46 3 62.83 93 75.15 11
SRL CoNLL-12 82.75 158 82.11 1 84.44 11 84.45 0 82.21 17 82.99 163
ID ATIS 97.98 2 98.09 0 98.04 1 97.98 0 97.70 1 97.48 1
SNIPS 98.43 0 98.43 0 96.72 0 96.72 0 98.00 0 98.15 0
DST MultiWOZ 39.61 0 39.35 0 41.02 0 41.02 0 38.85 518 39.86 4
Average 77.70 47.77 76.78 0.85 78.93 13.62 78.85 0.54 63.53 559.38 77.61 46.15
Table 10: Full result of F1 scores and format errors regarding to dataset specific fine-tuning.
Table  10 shows full result of fine-tuning (Section  6.1). Comparing with Table  9, fine-tuning shows 5 to 7 points better F1 score than formatted decoding. Surprisingly, best setting utilizing fine-tuning is using only format loss, which is better than using both formatted decoding. Model with format loss win over model with dataset information on almost every dataset. This is evidence that our format loss can reduce problem as easy as classification. However, using both fine-tuning and formatted decoding seems making conflict to each other. This is because some gold instances violates format restriction (e.g. <SOURCE> restriction in named entity recognition). Without formatted decoding, model can infer tokens that did not appeared in source sentence, when training example exists. However, we think this is not good for real-world usage, because it can be seen as overfitting on erroneous training examples.
C.3 Model size
Table 11: Full result regarding to model size. S, B, L stands for small, base, large, respectively. For each dataset, upper row is F1 score, and lower row is format errors. Text with bold means state of the art performance.
Task Dataset CE CE+FD CE+FL SLGM CE+Data SLGM+FT
S B L S B L S B L S B L S B L S B L
NER CoNLL-03 48.13 64.03 85.69 58.69 80.36 86.37 54.70 72.76 83.75 70.04 84.18 89.72 71.69 89.79 92.72 88.17 93.65 94.80
5684 5843 1756 9 43 5 6108 4437 2531 36 25 18 374 15 45 0 0 0
[…]
2080 717 555 360 191 42 1236 411 291 139 99 81 98 83 375 1 0 0
[…]
193 1137 5 8 15 7 197 31 16 12 7 4 92 35 0 1 0 0
[…]
7 5 9 8 2 2 56 7 9 4 2 4 4 7 14 4 2 2
[…]
7 5 9 8 2 2 56 7 9 4 2 4 7 7 15 4 2 2
[…]
690 794 643 0 0 0 702 874 635 0 0 0 420 11 1 0 0 0
0.00 4.87 45.77 7.98 27.28 45.07 0.00 23.32 38.61 9.32 33.90 44.84 0.00 46.43 59.69 26.00 55.24 60.02
470 530 151 91 0 7 883 425 208 244 2 2 981 99 19 1 0 1
[…]
564 475 93 322 232 939 446 215 77 200 95 47 16 37 0 0 0 0
45.68 54.79 84.66 45.20 66.74 84.98 29.99 65.01 85.22 29.63 64.75 85.21 49.72 66.01 86.06 38.95 76.46 86.66
298 82 3 5 1 2 600 120 8 59 11 1 68 38 5 45 3 0
[…]
167 20 161 2 1 0 83 14 9 4 0 1 164 16 162 4 0 1
[…]
134 31 4 17 0 0 123 8 3 5 0 0 96 10 5 0 0 0
[…]
6 1 1 0 0 0 2 0 0 0 0 0 8 0 0 0 0 0
[…]
106 10 4 0 0 0 36 0 0 0 0 0 155 4 4 0 0 4
[…]
800.46 742.31 261.08 63.85 37.46 77.38 809.85 503.77 292.00 54.38 18.69 12.46 191.00 27.85 49.62 4.62 0.54 0.77
Table  C.3 shows full result regarding to model size (Section  6.2). On cross entropy model, model can distinguish characteristics of dataset when model parameters increase. The power of formatted inference on score is reduced when parameter size increased, yet still effective reducing format errors. Comparing CE+FL with CE, using only format loss showed worse score and more format error on large sized model. However, when mixed with formatted decoding, it showed additional increase comparing with models using single methods. Additionally, when we use every method we tried in this paper (i.e. format loss, formatted decoding, using bigger model, fine-tuning on dataset), we achieved state of the art performance on CoNLL-03 named entity recognition task and ATIS intent detection task. This means our framework still has room for improvement.
Table 11: Full result regarding to model size. S, B, L stands for small, base, large, respectively. For each dataset, upper row is F1 score, and lower row is format errors. Text with bold means state of the art performance.